# On the Weight Spectrum of Rate-Compatible Polar Codes

### v2

@@ II-H Notations and Definitions
Let $\mathcal{C}(m,\bm{u}_{1}^{i})=\{\bm{c}|\bm{c}=\tilde{\bm{u}}_{1}^{N}\bm{F}_{N},\tilde{\bm{u}}_{1}^{i}=\bm{u}_{1}^{i}\}$ be the polar coset of $\bm{u}_{1}^{i}$. Define $\mathcal{C}_{P}(m,\bm{u}_{1}^{i},X)=\{\bm{c}_{[N]/X}|\bm{c}=\tilde{\bm{u}}_{1}^{N}\bm{F}_{N},\tilde{\bm{u}}_{1}^{i}=\bm{u}_{1}^{i}\}$ to be the polar coset of $\bm{u}_{1}^{i}$ under puncture pattern $X$. Note that if $X=[N]$, i.e., all bits are punctured, then there are $2^{m-i}$ empty codewords with length 0 and weight 0 in $\mathcal{C}_{P}(m,\bm{u}_{1}^{i},[N])$. Similarly, define $\mathcal{C}_{S}(m,\bm{u}_{1}^{i},Y)=\{\bm{c}_{[N]/Y}|\bm{c}=\tilde{\bm{u}}_{1}^{N}\bm{F}_{N},\tilde{\bm{u}}_{1}^{i}=\bm{u}_{1}^{i},\bm{c}_{Y}=\bm{0}\}$ to be the polar coset of $\bm{u}_{1}^{i}$ under shorten pattern $Y$. Note that when $Y$ is dominated [35], if $\exists j\leq i,j\in Y$, $\bm{u}_{j}=1$, then $C_{N}^{(i)}(\bm{u}_{1}^{i-1},\bm{u}_{i},Y)$ is empty since the codewords in coset contradict with the shorten pattern. Let $A_{d}(m,\bm{u}_{1}^{i},X)(A_{d}(m,\bm{u}_{1}^{i},Y))$ be the number of weight-$d$ codewords in $\mathcal{C}_{p}(m,\bm{u}_{1}^{i},X)(\mathcal{C}_{S}(m,\bm{u}_{1}^{i},Y))$.
@@ III Number of minimum-weight codewords in the bit-reversal shortened polar codes
In this section, we calculate the number of minimum-weight polynomials in the decreasing monomial code $C_{S}(\mathcal{I},Y_{i}^{\prime})$ with bit-reversal shortening pattern $Y_{i}^{\prime}$. Since shortening does not change the Hamming weight of the codewords, the number of minimum-weight codewords is exactly $\sum_{f\in\mathcal{I}_{r}}|T_{S}(f,Y_{i}^{\prime})|$.
We first propose a framework to calculate the number of minimum-weight codewords in general shortened polar codes. This analysis will also be applied to the calculation of Wang-Liu shortened polar codes.
From Theorem 1, the codewords in $T(f)$ can be represented as (1). When the monomial $g$ is shortened, a codeword $p$ survives if and only if the coefficient of $g$ in $p$ is 0. This imposes an equation on the relevant coefficients $a_{j,s}$. When consider multiple shortened monomials, a system of equations emerges. In general, it is challenging to solve such a complicated system. Notwithstanding, within the context of bit-reversal shortening, the equation system is indeed solvable.
@@
For bit-reversal shortening, the first shortened bit is the monomial $1$, so $p$ survives if and only if $a_{5,0}a_{3,0}=0$, i.e., polynomials with $a_{5,0}=a_{3,0}=1$ are no longer codewords in the shortened code. Therefore, $\lambda_{f}/4$ minimum-weight codewords are removed due to shortening, and the remaining codewords number is $43\lambda_{f}/$ 4.
Next, the monomial $x_{5}$ is shortened, which demands $a_{5,0}=0$, thus, polynomials with $a_{5,0}=1$ are no longer codewords. Since the case $a_{5,0}=a_{3,0}=1$ is already shortened, the newly removed polynomials are those with $a_{5,0}=1,a_{3,0}=0$. Thus, $\lambda_{f}/4$ more minimum-weight codewords are removed, and the remaining number of codewords with weight 8 in $T(f)$ is $\lambda_{f}/2$ after 2-bit bit-reversal shortening.
@@
Let $f,g\in\mathcal{M}$, ev$(f)$ be the $\bm{q}^{\prime}_{i}$-$th$ row of $\bm{F}_{N}$, ev$(g)$ be the $\bm{q}^{\prime}_{j}$-$th$ row of $\bm{F}_{N}$. If one of the following is true:
@@
Denote $f=x_{i_{1}}\cdots x_{i_{r}}=x_{1}^{a_{1}}\cdots x_{m}^{a_{m}}$ and $g=x_{j_{1}}\cdots x_{j_{t}}=x_{1}^{b_{1}}\cdots x_{m}^{b_{m}}$. If $g$ is a factor of $f$, $b_{i}=1$ implies $a_{i}=1$. Hence, $i=D(a_{m},\cdots,a_{1})\leq D(b_{m},\cdots,b_{1})=j$.
If $f\preccurlyeq g$ and deg$(f)=$ deg$(g)=r$, then $i_{s}\leq j_{s}$ for all $s\in\{1,\dots,r\}$. Therefore, $i=D(a_{m},\dots,a_{1})=\sum_{s=1}^{r}2^{i_{s}-1}+1<\sum_{s=1}^{r}2^{j_{s}-1}+1=D(b_{m},\cdots,b_{1})=j$. ∎
Next, we analyse the minimum-weight codewords that are removed exactly after the $i$-$th$ bit is shortened, i.e., the set $T_{S}(f,Y^{\prime}_{i-1})/T_{S}(f,Y^{\prime}_{i})$.
@@
Denote $f=x_{i_{1}}\cdots x_{i_{r}}=x_{1}^{a_{1}}\cdots x_{m}^{a_{m}}$ with $F_{f}=\{i_{1},...,i_{r}\}$ and $i_{1}<\dots<i_{r}$, $g=x_{j_{1}}\cdots x_{j_{t}}=x_{1}^{b_{1}}\cdots x_{m}^{b_{m}}$ with $F_{g}=\{j_{1},...,j_{t}\}$ and $j_{1}<\dots<j_{t}$. For bit-reversal shortening, suppose that the shortened bit $g$ is the $i$-$th$ shortened bit, i.e, $\bm{q}^{\prime}_{N-i}=D(b_{1},\dots,b_{m})$.
@@
If $g$ is a factor of $f$, the newly removed polynomials are
@@
with $a_{k,0}=1$ for all $k\in F_{f}/F_{g}$ and $a_{k,0}=0$ for all $k\in F_{g}$. Therefore, the number of newly removed polynomials in $T(f)$ is $\lambda_{f}/2^{r}$, i.e., $|T_{S}(f,Y^{\prime}_{i-1})/T_{S}(f,Y^{\prime}_{i})|=\lambda_{f}/2^{r}$.
@@
If $g\not\preccurlyeq f$, by Theorem 1, $g$ never appears in any polynomials in $T(f)$, so shortening $g$ will not influence the codewords in $T(f)$.
@@
Therefore, we only need to consider $g\preccurlyeq f$. It follows that $t\leq r$ and $i_{t-l}\leq j_{r-l}$ for all $l\in\{1,...,t\}$. Consider the following two cases:
@@
1) $g$ is not a factor of $f$, i.e. $F_{g}\not\subset F_{f}$. If $t=r$, according to Lemma 1, $f$ is shortened before $g$. Therefore, we only need to consider $t<r$. We are going to prove that $g$ does not appear in any polynomial $p\in T_{S}(f,Y^{\prime}_{i-1})$. Therefore, shortening $g$ does not remove polynomials in $T_{S}(f,Y^{\prime}_{i-1})$, i.e, $T_{S}(f,Y^{\prime}_{i-1})=T_{S}(f,Y^{\prime}_{i})$. Assume by contradiction, there exists some $p\in T_{S}(f,Y^{\prime}_{i-1})$ such that the coefficient of $g$ in $p$ is $1$. Since $g\preccurlyeq f$ and $t<r$, $g$ is shortened before $f$ if and only if there exists some $s\in\{1,...,t\}$ such that $i_{t-s}<j_{r-s}$ and $i_{t-l}=j_{r-l}$ for $l<s$. Then $F_{g}\cap B(f,t-s)=\varnothing$, so $a_{i_{t-s},k}x_{k}$ for $k\in B(f,t-s)$ do not appear in $g$. Therefore, the coefficient of $g$ must be the multiple of $a_{i_{t-s},0}$. Denote the coefficient of $x_{i_{t-s}}g$ to be $\tilde{a}$, then the coefficient of $g$ is $a_{i_{t-s},0}\tilde{a}$. Since $a_{i_{t-s},0}\tilde{a}=1$ implies $\tilde{a}=1$, the coefficient of $x_{i_{t-s}}g$ is 1. However, from Lemma 1, $x_{i_{t-s}}g$ has been shortened before $g$, which is a contradiction.
@@
2) $g$ is a factor of $f$, i.e., $F_{g}\subset F_{f}$. Then the coefficient of $g$ in $p$ is $\prod_{s\in F_{f}/F_{g}}a_{s,0}$. Define $\beta_{f}(i)$ to be the number of factors of $f$ in the last $i$ bits of the bit reversal sequence $\bm{q}^{\prime}$, which is the number of factors which have been shortened. Then $g$ is $\beta_{f}(i)$-$th$ shortened factors of $f$. We prove this part by induction on $\beta_{f}(i)$. When $\beta_{f}(i)=1$, i.e., $g=1$ and $F_{g}=\varnothing$, the shortened polynomials are those whose coefficient of monomial 1 is non-zero, i.e., $\prod_{s\in F_{f}}a_{s,0}=1$. Therefore, the shortened polynomials satisfies $a_{s,0}=1$ for all $s\in F_{f}$, and the total removed polynomials number is $\lambda_{f}/2^{r}$.
@@
For the inductive step $\beta_{f}(i)-1$ to $\beta_{f}(i)$, if $p\in T_{S}(f,Y^{\prime}_{i-1})/T_{S}(f,Y^{\prime}_{i})$, the coefficient of $g$ in $p$ is 1, which means $\prod_{s\in F_{f}/F_{g}}a_{s,0}=1$. From Lemma 1, any factor of $g$ has been already shortened before $g$. Each subset $S\subset F_{g}$ induces a factor $x_{1}^{a^{\prime}_{1}}\cdots x_{m}^{a^{\prime}_{m}}$ of $g$ with $S=\{1\leq k\leq m\mid a^{\prime}_{k}=1\}$. Then polynomials with $a_{s,0}=1$ for $s\in F_{f}/S$ and $a_{s,0}=0$ for $s\in S$ have been shortened. The only remaining polynomials shortened are the polynomials with $a_{j,0}=1$ for all $j\in F_{f}/F_{g}$ and $a_{j,0}=0$ for all $j\in F_{g}$. Therefore, the number of newly shortened polynomial is exactly $\lambda_{f}/2^{r}$.
@@
From Lemma 2, for a degree-$r$ polynomial $f$, when $i$-$th$ bit $g$ is shortened, the polynomials in $T(f)$ are removed from the code if and only if $g$ is a factor of $f$, and the number of removed polynomials is always $\lambda_{f}/2^{r}$.
For example, when $N=32$, $f=x_{3}x_{5}$, the polynomials in $T(f)$ are removed when monomials $1,x_{5},x_{3}$ are shortened, and the remaining number after each shortening step is $96,64,32$ respectively.
The number of minimum-weight codewords in bit-reversal shortened polar code immediately follows from Lemma 2.
@@
Suppose $C(\mathcal{I},Y_{i}^{\prime})$ is a shortened decreasing $r$-th order polar code with length $N=2^{m}$ and shortening pattern $Y_{i}^{\prime}$, where $Y_{i}^{\prime}$ is the last $i$ bits of the bit reversal sequence $\bm{q}^{\prime}$. Let $f$ be a monomial with degree $r$ and $d=2^{m-r}$ be the minimum weight, then $|T_{f}(d,Y_{i}^{\prime})|=\lambda_{f}(1-\beta_{f}(i)/2^{r})$. Here $\beta_{f}(i)$ is the number of factors of $f$ in $Y_{i}^{\prime}$. Then
@@
The Algorithm 1 concludes the procedure for calculating the number of minimum-weight codewords for bit-reversal shortened polar code.
Algorithm 1 Calculate the number of minimum-weight codewords in bit-reversal shortened polar code

### v3

@@ II-H Notations and Definitions
Let $\mathcal{C}(m,\bm{u}_{1}^{i})=\{\bm{c}|\bm{c}=\tilde{\bm{u}}_{1}^{N}\bm{F}_{N},\tilde{\bm{u}}_{1}^{i}=\bm{u}_{1}^{i}\}$ be the polar coset of $\bm{u}_{1}^{i}$. We define $\mathcal{C}_{P}(m,\bm{u}_{1}^{i},X)=\{\bm{c}_{[N]/X}|\bm{c}=\tilde{\bm{u}}_{1}^{N}\bm{F}_{N},\tilde{\bm{u}}_{1}^{i}=\bm{u}_{1}^{i}\}$ as the polar coset of $\bm{u}_{1}^{i}$ under puncture pattern $X$. Note that if $X=[N]$, i.e., all bits are punctured, then there are $2^{m-i}$ empty codewords with length 0 and weight 0 in $\mathcal{C}_{P}(m,\bm{u}_{1}^{i},[N])$. Similarly, we define $\mathcal{C}_{S}(m,\bm{u}_{1}^{i},Y)=\{\bm{c}_{[N]/Y}|\bm{c}=\tilde{\bm{u}}_{1}^{N}\bm{F}_{N},\tilde{\bm{u}}_{1}^{i}=\bm{u}_{1}^{i},\bm{c}_{Y}=\bm{0}\}$ as the polar coset of $\bm{u}_{1}^{i}$ under shorten pattern $Y$. Note that when $Y$ complies with binary domination [37], if there exists $j\leq i,j\in Y$ such that $\bm{u}_{j}=1$, then $C_{N}^{(i)}(\bm{u}_{1}^{i-1},\bm{u}_{i},Y)$ is empty since the codewords in the coset contradict the shorten pattern. Let $A_{d}(m,\bm{u}_{1}^{i},X)$ and $A_{d}(m,\bm{u}_{1}^{i},Y)$ be the number of weight-$d$ codewords in $\mathcal{C}_{p}(m,\bm{u}_{1}^{i},X)$ and $\mathcal{C}_{S}(m,\bm{u}_{1}^{i},Y)$, respectively.
@@ III Number of minimum-weight codewords in the bit-reversal shortened polar codes
In this section, we calculate the number of minimum-weight polynomials in the $r$-thth order decreasing monomial code $C_{S}(\mathcal{I},Y_{i}^{\prime})$ with bit-reversal shortening pattern $Y_{i}^{\prime}$. Since shortening does not change the Hamming weight of the codewords, the number of minimum-weight codewords is exactly $\sum_{f\in\mathcal{I}_{r}}|T_{S}(f,Y_{i}^{\prime})|$.
We first propose a framework to calculate the number of minimum-weight codewords in general shortened polar codes. This analysis will also be applied to Wang-Liu shortened polar codes.
From Theorem 1, the codewords in $T(f)$ can be represented as (1). When the monomial $g$ is shortened, a codeword $p$ is valid if and only if the coefficient of $g$ in $p$ is 0. This imposes an equation on the relevant coefficients $a_{j,s}$. When considering multiple shortened monomials, a system of equations emerges. In general, it is challenging to solve such a complicated system. Nevertheless, within the context of bit-reversal shortening, the equation system is indeed solvable.
@@
For bit-reversal shortening, the first shortened bit corresponds to the monomial $1$, so $p$ is valid if and only if $a_{5,0}a_{3,0}=0$, i.e., polynomials with $a_{5,0}=a_{3,0}=1$ are excluded from the code. Therefore, $\lambda_{f}/4$ minimum-weight codewords are removed due to shortening, and the number of remaining codewords is $3\lambda_{f}/4$.
Next, the monomial $x_{5}$ is shortened, which requires $a_{5,0}=0$. Consequently, polynomials with $a_{5,0}=1$ are excluded from the code. Since the case $a_{5,0}=a_{3,0}=1$ is already shortened, the newly removed polynomials are those with $a_{5,0}=1,a_{3,0}=0$. Therefore, $\lambda_{f}/4$ more minimum-weight codewords are removed, and the number of remaining codewords with minimum weight in $T(f)$ is $\lambda_{f}/2$ after 2-bit bit-reversal shortening.
@@
Let $f,g\in\mathcal{M}$, ev(f)(f) be the $q^{\prime}_{i}$-thth row of $\bm{F}_{N}$, ev(g)(g) be the $q^{\prime}_{j}$-thth row of $\bm{F}_{N}$. If one of the following is true:
@@
Denote $f=x_{i_{1}}\cdots x_{i_{r}}=x_{1}^{a_{1}}\cdots x_{m}^{a_{m}}$ and $g=x_{j_{1}}\cdots x_{j_{t}}=x_{1}^{b_{1}}\cdots x_{m}^{b_{m}}$. If $g$ is a factor of $f$, then $b_{i}=1$ implies $a_{i}=1$. Hence, $i=D(a_{m},\cdots,a_{1})\leq D(b_{m},\cdots,b_{1})=j$.
If $f\preccurlyeq g$ and deg(f)=(f)= deg(g)=r(g)=r, then $i_{s}\leq j_{s}$ for all $s\in\{1,\dots,r\}$. Therefore, $i=D(a_{m},\dots,a_{1})=\sum_{s=1}^{r}2^{i_{s}-1}+1<\sum_{s=1}^{r}2^{j_{s}-1}+1=D(b_{m},\cdots,b_{1})=j$. ∎
Next, we analyse the minimum-weight codewords that are removed exactly after the $i$-thth bit is shortened, i.e., the set $T_{S}(f,Y^{\prime}_{i-1})/T_{S}(f,Y^{\prime}_{i})$.
@@
Let $f=x_{i_{1}}\cdots x_{i_{r}}=x_{1}^{a_{1}}\cdots x_{m}^{a_{m}}$ with $F_{f}=\{i_{1},...,i_{r}\}$ and $i_{1}<\dots<i_{r}$, $g=x_{j_{1}}\cdots x_{j_{t}}=x_{1}^{b_{1}}\cdots x_{m}^{b_{m}}$ with $F_{g}=\{j_{1},...,j_{t}\}$ and $j_{1}<\dots<j_{t}$. For bit-reversal shortening, suppose that $g$ is the $i$-thth shortened monomial.
@@
If $g$ is a factor of $f$, the newly removed polynomials are those of the form
@@
with $a_{k,0}=1$ for all $k\in F_{f}/F_{g}$ and $a_{k,0}=0$ for all $k\in F_{g}$. Therefore, the number of newly removed polynomials from $T(f)$ is $\lambda_{f}/2^{r}$, which means $|T_{S}(f,Y^{\prime}_{i-1})/T_{S}(f,Y^{\prime}_{i})|=\lambda_{f}/2^{r}$.
@@
If $g\not\preccurlyeq f$, by Theorem 1, $g$ never appears in any polynomial from $T(f)$, so shortening $g$ will not influence the polynomials in $T(f)$.
@@
Therefore, we only need to consider $g\preccurlyeq f$, which means $t\leq r$ and $i_{t-l}\leq j_{r-l}$ for all $l\in\{1,...,t\}$. Consider the following two cases:
@@
1) $g$ is not a factor of $f$, i.e., $F_{g}\not\subset F_{f}$. If $t=r$, according to Lemma 1, $f$ is shortened before $g$. Therefore, we only need to consider the case where $t<r$. We will prove that $g$ does not appear in any polynomial $p\in T_{S}(f,Y^{\prime}_{i-1})$. Thus, shortening $g$ does not remove polynomials from $T_{S}(f,Y^{\prime}_{i-1})$, i.e., $T_{S}(f,Y^{\prime}_{i-1})=T_{S}(f,Y^{\prime}_{i})$.
Now, assume for contradiction that there exists some $p\in T_{S}(f,Y^{\prime}_{i-1})$ such that the coefficient of $g$ in $p$ is $1$. Since $g\preccurlyeq f$ and $t<r$, $g$ is shortened before $f$ if and only if there exists some $s\in\{1,...,t\}$ such that $i_{t-s}<j_{r-s}$ and $i_{t-l}=j_{r-l}$ for $l<s$. Then $F_{g}\cap B(f,t-s)=\varnothing$, so $a_{i_{t-s},k}x_{k}$ for $k\in B(f,t-s)$ do not appear in $g$. Therefore, the coefficient of $g$ must be a multiple of $a_{i_{t-s},0}$. Denote the coefficient of $x_{i_{t-s}}g$ to be $\tilde{a}$, then the coefficient of $g$ is $a_{i_{t-s},0}\tilde{a}$. Since $a_{i_{t-s},0}\tilde{a}=1$ implies $\tilde{a}=1$, the coefficient of $x_{i_{t-s}}g$ is 1. However, according to Lemma 1, $x_{i_{t-s}}g$ has been shortened before $g$, which is a contradiction.
@@
2) $g$ is a factor of $f$, which means $F_{g}\subset F_{f}$. Then the coefficient of $g$ in $p$ is $\prod_{s\in F_{f}/F_{g}}a_{s,0}$. Define $\beta_{f}(i)$ to be the number of factors of $f$ in the last $i$ bits of the bit reversal sequence $\bm{q}^{\prime}$, which is the number of factors that have been shortened. Then $g$ is $\beta_{f}(i)$-thth shortened factor of $f$. We prove this by induction on $\beta_{f}(i)$. When $\beta_{f}(i)=1$, i.e., $g=1$ and $F_{g}=\varnothing$, the shortened polynomials are those whose coefficient of monomial 1 is non-zero, i.e., $\prod_{s\in F_{f}}a_{s,0}=1$. Therefore, the shortened polynomials must satisfy $a_{s,0}=1$ for all $s\in F_{f}$, and the total number of removed polynomials is $\lambda_{f}/2^{r}$.
@@
For the inductive step $\beta_{f}(i)-1$ to $\beta_{f}(i)$, if $p\in T_{S}(f,Y^{\prime}_{i-1})/T_{S}(f,Y^{\prime}_{i})$, the coefficient of $g$ in $p$ is 1, which means $\prod_{s\in F_{f}/F_{g}}a_{s,0}=1$. From Lemma 1, any factor of $g$ has been shortened before $g$. Let $x_{1}^{a^{\prime}_{1}}\cdots x_{m}^{a^{\prime}_{m}}$ of $g$ with $S=\{1\leq k\leq m\mid a^{\prime}_{k}=1\}\subset F_{g}$ be a factor of $g$, and of course, a factor of $f$, whose coefficient in $p$ is $\prod_{s\in F_{f}/S}a_{s,0}$. Thus, polynomials with $a_{s,0}=1$ for $s\in F_{f}/S$ and $a_{s,0}=0$ for $s\in S$ have been shortened. The only remaining polynomials to be shortened are those with $a_{j,0}=1$ for all $j\in F_{f}/F_{g}$ and $a_{j,0}=0$ for all $j\in F_{g}$. Therefore, the number of newly shortened polynomials is exactly $\lambda_{f}/2^{r}$.
@@
From Lemma 2, for a degree-$r$ polynomial $f$, when the monomial $g$ is shortened, the polynomials in $T(f)$ are removed from the code if and only if $g$ is a factor of $f$, and the number of removed polynomials is always $\lambda_{f}/2^{r}$.
For example, when $N=32$, $f=x_{3}x_{5}$, the polynomials in $T(f)$ are removed when monomials $1,x_{5},x_{3}$ are shortened, and the remaining number of polynomials after each shortening step is $96,64,32$, respectively.
The number of minimum-weight codewords in bit-reversal shortened polar codes immediately follows from Lemma 2.
@@
Let $C(\mathcal{I},Y_{i}^{\prime})$ be a shortened decreasing $r$-th order polar code with length $N=2^{m}$ and shortening pattern $Y_{i}^{\prime}$, where $Y_{i}^{\prime}$ is the last $i$ bits of the bit reversal sequence $\bm{q}^{\prime}$. Let $f$ be a monomial with degree $r$ and $d=2^{m-r}$ be the minimum weight, then $|T_{f}(d,Y_{i}^{\prime})|=\lambda_{f}(1-\beta_{f}(i)/2^{r})$. Here $\beta_{f}(i)$ is the number of factors of $f$ in $Y_{i}^{\prime}$. Then
@@
The Algorithm 1 describes the procedure for calculating the number of minimum-weight codewords for bit-reversal shortened polar codes.
Algorithm 1 Calculate the number of minimum-weight codewords in bit-reversal shortened polar codes